OpenStudy (anonymous):

trig help! solve each equation for solutions in the interval [0,2π) by first solving for the trigonometric function. sec^2 x + 2 = -1 OR (cotx - 1)(√3 cotx + 1) = 0 i'm stuck on both.. please help. any explanations are welcome..

OpenStudy (anonymous):

for the second question, i set each to 0 and i got cotx = 1 and cot x = -1/√3 but i'm not sure what the next step is... the answers in the back of the book are π/4, 2π/3, 5π/4, and 5π/3

OpenStudy (callisto):

Consider\[\cot x = 1\]Multiply both sides by tanx\[1=\tan x\]What value of x would tanx be 1?

OpenStudy (anonymous):

π/4?

OpenStudy (callisto):

Yes! One more

OpenStudy (anonymous):

that's the problem.. i only know how to find the first answer. i don't know how to find the second one :( do i add 2π?

OpenStudy (callisto):

Note than tanx = tan(π+ x)

OpenStudy (anonymous):

so π/4 + 2π? 9π/4? i'm sorry.. i don't know what you mean..

OpenStudy (callisto):

No :( Do you know that tan(π+x) = tanx?

OpenStudy (anonymous):

no.. wait i think so. do you mean that the two angles are the same? like 45 and 225 degrees?

OpenStudy (callisto):

The two angles are NOT the same, BUT tan(45) = tan(225) :)

OpenStudy (anonymous):

oh. then what do you mean by tan(π+x) = tanx? i've never heard of it..

OpenStudy (callisto):

pi = 180 degrees. Let say, we have x = 45 tan(45) = tan(180+45) = 1 tan(180+45) is the same as tan(225). Got it?

OpenStudy (anonymous):

OH. yes i get it :D so essentially the two angles are the same right? so i could add 180 to 45 as many times as i want and the two angles would still be alike?

OpenStudy (callisto):

For tan, yes. That is tan(x) = tan(nπ +x)

OpenStudy (anonymous):

okay thank you! :) i think i know what i've been doing wrong. to find other angles, i've been adding 360(2π) instead of 180(π). so back to my question.. i would solve for cotx and graph it. the first answer is π/4, and i would add π to get the second answer, right?

OpenStudy (callisto):

Yes.

OpenStudy (callisto):

So, the answers to (cot-1)=0 are...??

OpenStudy (anonymous):

π/4 and 5π/4?

OpenStudy (callisto):

Yes

OpenStudy (anonymous):

thanks so much ! :D

OpenStudy (callisto):

Now, can you solve √3 cotx + 1 =0?

OpenStudy (anonymous):

√3 cotx + 1 =0 √3 cotx = -1 cotx = -1/√3 did i draw it right?